OpenStudy (anonymous):

Find the vertex of y=x^2+4x-5. show your work.

OpenStudy (anonymous):

1) take the derivative \[y = x^{2} + 4x - 5\] => \[y' = 2x + 4\] 2) set the derivative equal to zero and solve for x \[2x+4=0 => x=-2\] 3) plug the x you found back into the original equation \[y = (-2)^2 + 4(-2) - 5 = -9\] 4) you have both the x and y coordinate now.

OpenStudy (anonymous):

how did you get from y=x^2+4x-5 to y=2x+4?

OpenStudy (anonymous):

@chrismoon

OpenStudy (anonymous):

It's y', not y. Have you learned about derivatives yet?

OpenStudy (anonymous):

no.. i'm confused.

OpenStudy (anonymous):

Check this page out. Should help: http://www.purplemath.com/modules/sqrvertx.htm

OpenStudy (anonymous):

tthanks.

OpenStudy (anonymous):

you could also complete the square. longer route, but good for algebra students :)